# Paralinguistics-Enhanced Large Language Modeling of Spoken Dialogue

### 23 Dec 2023 v1

@@ Abstract
Large Language Models (LLMs) have demonstrated superior abilities in tasks such as chatting, reasoning, and question-answering. However, standard LLMs may ignore crucial paralinguistic information, such as sentiment, emotion, and speaking style, which are essential for achieving natural, human-like spoken conversation, especially when such information is conveyed by acoustic cues. We therefore propose Paralinguistics-enhanced Generative Pretrained Transformer (ParalinGPT), an LLM utilizes text and speech modality to better model the linguistic content and paralinguistic attribute of spoken response. The model takes the conversational context of text, speech embeddings, and paralinguistic attributes as input prompts within a serialized multitasking multi-modal framework. Specifically, our framework serializes tasks in the order of current paralinguistic attribute prediction, response paralinguistic attribute prediction, and response text generation with autoregressive conditioning. We utilize the Switchboard-1 corpus, including its sentiment labels to be the paralinguistic attribute, as our spoken dialogue dataset. Experimental results indicate the proposed serialized multitasking method outperforms typical sequence classification techniques on current and response sentiment classification. Furthermore, leveraging conversational context and speech embeddings significantly improves both response text generation and sentiment prediction. Our proposed framework achieves relative improvements of 6.7%, 12.0%, and 3.5% in current sentiment accuracy, response sentiment accuracy, and response text BLEU score, respectively.
@@ 1 Introduction
Despite the existence of prior work that leverages sentiment or emotion labels to enhance dialogue modeling [4, 5, 6, 7, 8, 9, 2], these studies are primarily based on textual dialogue data, with sentiment or emotion labels being inferred from the word alone. On the other hand, prior work in modeling spoken dialogue has also shown that prosodic information is helpful for language modeling, i.e., predicting the words [10, 11].
In this work, alongside the linguistic information extracted from text, we introduce ParalinGPT, a language model capable of incorporating speech embeddings from a self-supervised speech encoder as input prompts, within a serialized multitasking multi-modal framework designed to predict upcoming text- and speech-based paralinguistic information. We employ the Switchboard-1 dataset, the largest collection of human-human spontaneous spoken dialogues with annotated speech sentiments, as our primary spoken dialogue data source.
@@
To our knowledge, this is a novel attempt to employ a multi-modal language model for multiple tasks involving both response text generation and prediction of response and current sentiment labels, particularly within the context of real-world spoken dialogues. Our specific contributions include:
@@
  • Results showing that the framework achieves superior performance compared to sequence classification baselines, with 6.7% and 12.0% accuracy improvement for current and response sentiment, respectively, while seeing 3.5% relative improvement in BLEUd score for response text generation.
@@ 2.1 Dialogue Modeling with Sentiment and Emotion
There are several prior works aiming to model the sentiment [4, 5, 12, 13, 14, 15] or emotion [12, 6, 7, 8, 9] in dialogue. Most of this work concentrates on current-turn sentiment/emotion as a pure classification task. However, previous studies have also shown that using multi-modal approaches are crucial for sentiment and emotion analysis [16, 17, 18, 19, 20, 21].
There have been fewer attempts at response modeling where the goal is to predict the response sentiment [4] or emotion [12] label; others are leveraging the sentiment [15, 5] or emotion label [6, 7, 8, 9, 2] to guide the response text generation. It is worth noting that Varshney et al. [8] proposed a multi-tasking framework with context-turn emotion recognition and response generation, showing improved emotional relevancy of the generated response, but they do not predict the response emotion. Compared to this prior work, our proposal is novel in the following aspects: (i) we use multi-modal input (speech and text) with multi-turn dialog context on real-world spoken dialogue, and (ii) we model current and response sentiment, as well as response text as conditioned on the paralinguistic labels.
@@ 2.2 Language Modeling for Speech
Much recent research uses spoken language modeling on discrete speech/audio units [22, 23, 24, 25, 26]. [22, 23, 24, 25] use a decoder-only LM to perform speech continuation. AudioPaLM [26] and SpeechGPT [27] leverage paired, aligned transcripts and speech for ASR and TTS tasks. However, discrete audio/speech units mainly capture content information and lack prosodic and paralinguistic information. pGSLM [23] explicitly models extra pitch and duration units to improve speech naturalness. Notably, dGSLM [24] is the only study of two-channel language modeling for spoken dialogue, but the generated speech is still lacking in semantic coherence. Different from the above-mentioned studies, we focus on equipping a strong text LM with continuous speech embeddings to capture the paralinguistic signal in spoken dialogue.
@@ 3.1 The Serialized Multitasking Framework
Inspired by the above and to move beyond isolated predictions, we propose a sequential conditioning mechanism, encompassing current sentiment prediction, response sentiment prediction, and response text generation, all joined in an autoregressive chain; we call this framework serialized multitasking since the model’s input and output streams encode multiple tasks. The proposed framework is illustrated in Figure 2.
We formulate the problem as follows: given a multi-turn spoken dialog with historical context turns $H$ and window size $w$, comprising speech $\{S_{n-w},..,S_{n-1}\}$, text $\{T_{n-w},..,T_{n-1}\}$, and sentiment labels $\{L_{n-w},..,L_{n-1}\}$, a current turn $C$ comprising speech $S_{n}$ and $T_{n}$, our prediction target $O$ is the current sentiment $L_{n}$, response sentiment $L_{n+1}$, and response text $T_{n+1}$. The overall objective is as follows:
@@
The training objective is the typical cross-entropy loss for language modeling with teacher forcing. During testing, we generate $O$ by autoregressively feeding earlier predicted tokens back into the LLM inputs to predict a target sequence that contains current sentiment, response sentiment, and response text all presented as text. Note that the text and sentiment label in the history $H$, as well as the current text tokens $T_{n}$ are ground truth. Future work can look at using ASR output and erroneous versions of historical sentiment labels. After sequence generation, we use the special marker <> to decode sentiment predictions for evaluation.
@@ 3.2 Language Model
Method # Current Context History Prediction Target Modality Modality Sentiment Curr UA (normal-$\uparrow$) Res UA (normal-$\uparrow$) BLEU (normal-$\uparrow$) Lu et al. [29] 1 E2E ASR 62.4 - - Shon et al. [30] 2 T 64.5 - - Human performance [29] 3 84.6 - - Random baseline 4 33.3 33.3 - LM only 5 T - - 14.2 6 T T - - 14.5 Sequence Classification $*{}^{*}$ 7 T 66.3 36.1 - 8 S 65.1 34.8 - 9 T + S 68.1 37.4 - Proposed Serialized Multitasking Framework 10 T 68.3 38.0 14.0 11 T T 69.3 40.5 14.4 12 T T 70.3 41.4 14.5 13 T + S 70.7 38.6 14.1 14 T + S T + S 71.3 40.9 14.7 15 T + S T + S 71.7 41.9 14.9
Table 1: Experimental results: “Curr” and “Res” denote current and response sentiment, respectively. In the “Modality” column, “T” indicates text, and “S” denotes speech. “✓” indicates presence and “✗” indicates absence. $*{}^{*}$s a sequence classification baseline, we train separate models for current sentiment and response sentiment.
@@ 4.2 Implementation Details
We utilize the pretrained DialoGPT with 345M parameters [^1] https://huggingface.co/microsoft/DialoGPT-medium as LM and wav2vec 2.0 Large as the speech encoder [^2] https://huggingface.co/facebook/wav2vec2-large-robust. For the context history, we use window size $w=4$ and truncate the input if the length exceeds 320 tokens during training. We use a batch size of 32 and Adam optimizer with learning rate 0.0001, and train for 100k steps. During training, the parameters of the speech encoder are frozen and only the parameters of the LM and speech linear projector are updated. We select the best model according to the validation set, and report the performance on the test set.
@@ 4.3 Baseline approaches
We consider several baselines to benchmark our proposed model for each task, based on prior published results.
Response text generation: We fine-tune DialoGPT on Switchboard data using typical LM objective and multi-turn context history as a baseline.
Sentiment classification: We use three baseline systems to benchmark sentiment classification against our proposed technique. (i) Random baseline, since ours is the first work targeting response sentiment prediction on the Switchboard corpus. (ii) Sequence classification: for speech, we fine-tune wav2vec-2.0 with a classifier head; for text, we fine-tune DialoGPT for simple classification; for multi-modal approaches combining text and speech we augment DialoGPT with input text concatenated with wav2vec 2.0 mean-pooled embeddings. (iii) Prior published results on the Switchboard corpus for utterance classification.
@@ 5.1 Baselines
From prior work on current-utterance sentiment analysis, we quote the performance reported in [29, 30], while noting differences in terms of the data split. For response sentiment, we utilize a random (chance) baseline, since there is no prior work on Switchboard to compare to. The sequence classification baseline is used to assess the task difficulty and for comparison to our proposed serialized multitasking method.
@@ 5.2 Serialized multitasking framework
5.3 Multi-modal serialized multitasking framework
@@
Task order Curr UA Res UA BLEU $L_{n}\rightarrow L_{n+1}\rightarrow T_{n+1}$ (Proposed) 70.3 41.4 14.5 $L_{n}\rightarrow T_{n+1}\rightarrow L_{n+1}$ 64.6 38.8 13.1 $L_{n+1}\rightarrow T_{n+1}$ 40.9 14.4
@@ 5.4 Role of Multi-turn Context
Multi-modal context: Beyond text context, we can incorporate speech embeddings for contextual turns to provide more acoustic clues for interpretation and prediction. Comparing text (Table 1, row 11) with Text+Speech context (row 14), we observe further improvements of 1.4%, 0.5%, and 2.1% in relative terms for current sentiment, response sentiment, and BLEU score, respectively.
Sentiment context: Lastly, using contextual sentiment labels also enhances the overall performance of the proposed method. For the text modality, there are relative gains of 1.4%, 1.7%, and 0.7% in current sentiment, response sentiment, and response text, respectively. Overall, the best-performing multi-modal serialized method (row 15) achieves a current sentiment accuracy of 71.7%, response sentiment accuracy of 41.9%, and a response text BLEU score of 14.9. This result implies that contextual speech embeddings and sentiment information are complementary. We hypothesize that the rich information in speech, such as speaker identity, helps to disambiguate and understand the spoken dialogue effectively.
@@ 6 Conclusion
We introduced ParalinGPT, a modeling framework for spoken dialogues that integrates paralinguistic and linguistic information in a single autoregressive model. This framework enables the joint generation of response text and current and response sentiment prediction. Through our experiments on a corpus of human-human conversations, we show the effectiveness of the proposed serialized multitasking method compared to separate classification or language models. Additionally, we demonstrate that incorporating self-supervised speech features and context history further improves performance. There are several potential avenues for future research, including using the latest and large-scale language models and adding more tasks under serialized multi-tasking, such as turn taking prediction and emotion recognition. Another important direction for future research involves the exploration of paralinguistic speech embeddings at various time resolutions and pooling methods.

### 17 Jan 2024 v2

@@ Abstract
Large Language Models (LLMs) have demonstrated superior abilities in tasks such as chatting, reasoning, and question-answering. However, standard LLMs may ignore crucial paralinguistic information, such as sentiment, emotion, and speaking style, which are essential for achieving natural, human-like spoken conversation, especially when such information is conveyed by acoustic cues. We therefore propose Paralinguistics-enhanced Generative Pretrained Transformer (ParalinGPT), an LLM that utilizes text and speech modalities to better model the linguistic content and paralinguistic attributes of spoken dialogue. The model takes the conversational context of text, speech embeddings, and paralinguistic attributes as input prompts within a serialized multitasking multimodal framework. Specifically, our framework serializes tasks in the order of current paralinguistic attribute prediction, response paralinguistic attribute prediction, and response text generation with autoregressive conditioning. We utilize the Switchboard-1 corpus, including its sentiment labels as the paralinguistic attribute, as our spoken dialogue dataset. Experimental results indicate the proposed serialized multitasking method outperforms typical sequence classification techniques on current and response sentiment classification. Furthermore, leveraging conversational context and speech embeddings significantly improves both response text generation and sentiment prediction. Our proposed framework achieves relative improvements of 6.7%, 12.0%, and 3.5% in current sentiment accuracy, response sentiment accuracy, and response text BLEU score, respectively.
@@ 1 Introduction
Despite prior work that leverages sentiment or emotion labels to enhance dialogue modeling [4, 5, 6, 7, 8, 9, 2], such studies are primarily based on textual dialogue data, with sentiment or emotion labels being inferred from the word alone. On the other hand, prior work in modeling spoken dialogue has also shown that prosodic information is helpful for language modeling, i.e., predicting the words [10, 11].
In this work, moving beyond information extracted from text alone, we introduce ParalinGPT, a language model capable of incorporating speech embeddings from a self-supervised speech encoder as input prompts, within a serialized multitasking multimodal framework designed to predict both upcoming text- and speech-based paralinguistic information. We employ the Switchboard-1 dataset, the largest collection of human-human spontaneous spoken dialogues with annotated “speech sentiments”, as our primary spoken dialogue data source.
@@
To our knowledge, this is the first attempt to employ a multimodal language model for joint response text generation and prediction of response and current sentiment labels, particularly in the domain of real-world spoken dialogues. Our specific contributions include:
@@
  • Results showing that the framework achieves superior performance compared to sequence classification baselines, with 6.7% and 12.0% accuracy improvement for current and response sentiment, respectively, while seeing 3.5% relative improvement in BLEU score for response text generation.
@@ 2.1 Dialogue Modeling with Sentiment and Emotion
There are several prior studies aiming to model sentiment [4, 5, 12, 13, 14, 15] and emotion [12, 6, 7, 8, 9] in dialogue. Most of this work concentrates on current-turn sentiment/emotion as a pure classification task. However, previous work has also shown that using multimodal approaches is crucial for sentiment and emotion analysis [16, 17, 18, 19, 20, 21].
There have been fewer attempts at response modeling where the goal is to predict the response sentiment [4] or emotion [12, 22, 23] label; others are leveraging the sentiment [15, 5] or emotion label [6, 7, 8, 9, 2] to guide the response text generation. It is worth noting that Varshney et al. [8] proposed a multitask framework with context-turn emotion recognition and response generation, showing improved emotional relevancy of the generated response, but they do not predict the response emotion. Compared to this prior work, our proposal is novel in the following aspects: (i) we use multimodal input (speech and text) with multi-turn dialog context on real-world spoken dialogue, and (ii) we model current and response sentiment, as well as response text as conditioned on the paralinguistic labels.
@@ 2.2 Language Modeling for Speech
Much recent research uses spoken language modeling on discrete speech/audio units, either using a decoder-only LM to perform speech continuation [24, 25, 26, 27] or leveraging paired transcripts and speech for ASR and TTS tasks [28, 29]. However, discrete audio/speech units mainly capture content information and lack prosodic and paralinguistic information. pGSLM [25] explicitly models extra pitch and duration units to improve speech naturalness. Notably, dGSLM [26] is the only study of two-channel language modeling for spoken dialogue, but the generated speech is still lacking in semantic coherence. Different from the above-mentioned studies, we focus on equipping a strong text LM with continuous speech embeddings to capture the paralinguistic signal in spoken dialogue.
@@ 3.1 The Serialized Multitasking Framework
Inspired by the above rationale, and to move beyond isolated predictions, we propose a sequential conditioning mechanism, encompassing current sentiment prediction, response sentiment prediction, and response text generation, all joined in an autoregressive chain; we call this framework serialized multitasking since the model’s input and output streams encode multiple tasks. The proposed framework is illustrated in Figure 2.
We formalize the problem as follows: given a multi-turn spoken dialog with historical context turns $H$ and window size $w$, comprising speech $\{S_{n-w},..,S_{n-1}\}$, text $\{T_{n-w},..,T_{n-1}\}$, and sentiment labels $\{L_{n-w},..,L_{n-1}\}$, a current turn $C$ comprising speech $S_{n}$ and text $T_{n}$, our prediction target $O$ is the current sentiment $L_{n}$, response sentiment $L_{n+1}$, and response text $T_{n+1}$. The overall objective is as follows:
@@
The training objective is the typical cross-entropy loss for language modeling with teacher forcing. During testing, we generate $O$ by autoregressively feeding earlier predicted tokens back into the LLM inputs to predict a target sequence that contains current sentiment, response sentiment, and response text all represented as text. Note that the text and sentiment label in the history $H$, as well as the current text tokens $T_{n}$ are ground truth. Future work can look at using ASR output and erroneous versions of historical sentiment labels. After sequence generation, we use the special marker <> to decode sentiment predictions for evaluation.
@@ 3.2 Language Model
Table 1: Experimental results: “Curr” and “Res” denote current and response sentiment, respectively. In the “Modality” column, “T” indicates text, and “S” denotes speech. “✓” indicates presence and “✗” indicates absence of context information. As sequence classification baselines (${}^{\ast}$) we train separate models for current sentiment (Curr UA) and response sentiment (Resp UA).
Method # Current Context History Prediction Target Modality Modality Sentiment Curr UA (normal-$\uparrow$) Resp UA (normal-$\uparrow$) BLEU (normal-$\uparrow$) Lu et al. [31] 1 E2E ASR 62.4 - - Shon et al. [32] 2 T 64.5 - - Human performance [31] 3 84.6 - - Random baseline 4 33.3 33.3 - LM only 5 T - - 14.2 6 T T - - 14.5 Sequence Classification $*{}^{*}$ 7 T 66.3 36.1 - 8 S 65.1 34.8 - 9 T + S 68.1 37.4 - Proposed Serialized Multitasking Framework 10 T 68.3 38.0 14.0 11 T T 69.3 40.5 14.4 12 T T 70.3 41.4 14.5 13 T + S 70.7 38.6 14.1 14 T + S T + S 71.3 40.9 14.7 15 T + S T + S 71.7 41.9 14.9
@@ 4.2 Implementation Details
We utilize the pretrained DialoGPT with 345M parameters [38] as LM and wav2vec 2.0 Large as the speech encoder [39]. For the context history, we use window size $w=4$ and truncate the input if the length exceeds 320 tokens during training. We use a batch size of 32 and Adam optimizer with learning rate 0.0001, and train for 100k steps. During training, the parameters of the speech encoder are frozen and only the parameters of the LM and speech linear projector are updated. We select the best model according to the validation set, and report the performance on the test set.
@@ 4.3 Baseline Approaches
For each task, we consider several baselines against which to benchmark our proposed model, based on published results.
Response text generation: We fine-tune DialoGPT on Switchboard data using the standard LM objective and multi-turn context history as a baseline.
Sentiment classification: We use three baseline systems to benchmark sentiment classification against our proposed technique. (i) Random baseline, since ours is the first work targeting response sentiment prediction on the Switchboard corpus. (ii) Sequence classification: for speech, we fine-tune wav2vec-2.0 with a classifier head; for text, we fine-tune DialoGPT for simple classification; for multimodal approaches combining text and speech we augment DialoGPT with input text concatenated with wav2vec 2.0 mean-pooled embeddings. (iii) Prior published results on the Switchboard corpus for utterance classification.
@@ 5.1 Baselines
From prior work on single-utterance sentiment analysis, we quote the performance reported in [31, 32], while noting differences in terms of the data split. For response sentiment, we utilize a random (chance) baseline, since there is no prior work on Switchboard to compare to. The sequence classification baseline is used to assess the task difficulty and for comparison to our proposed serialized multitasking method.
@@ 5.2 Serialized multitasking framework
5.3 Multimodal serialized multitasking framework
@@
Task Order Curr UA Resp UA BLEU $L_{n}\rightarrow L_{n+1}\rightarrow T_{n+1}$ (Proposed) 70.3 41.4 14.5 $L_{n}\rightarrow T_{n+1}\rightarrow L_{n+1}$ 64.6 38.8 13.1 $L_{n+1}\rightarrow T_{n+1}$ 40.9 14.4
@@ 5.4 Role of Multi-turn Context
Multimodal context: Beyond text context, we can incorporate speech embeddings for contextual turns to provide more acoustic clues for interpretation and prediction. Comparing text (Table 1, row 11) with Text+Speech context (row 14), we observe further improvements of 1.4%, 0.5%, and 2.1% in relative terms for current sentiment, response sentiment, and BLEU score, respectively.
Sentiment context: Lastly, using contextual sentiment labels also enhances the overall performance of the proposed method. For the text modality, there are relative gains of 1.4%, 1.7%, and 0.7% in current sentiment, response sentiment, and response text, respectively. Overall, the best-performing multimodal serialized method (row 15) achieves a current sentiment accuracy of 71.7%, response sentiment accuracy of 41.9%, and a response text BLEU score of 14.9. This result implies that contextual speech embeddings and sentiment information are complementary. We hypothesize that the rich information in speech, such as timing and intonation, helps to disambiguate and understand the spoken dialogue more effectively.
@@ 6 Conclusion
We have introduced ParalinGPT, a modeling framework for spoken dialogues that integrates paralinguistic and linguistic information in a single autoregressive model. This framework enables the joint generation of response text and current and response sentiment prediction. Through our experiments on a corpus of human-human conversations, we show the effectiveness of the proposed serialized multitasking method compared to separate classification and language models. Additionally, we demonstrate that incorporating self-supervised speech features and context history further improves performance. There are several directions for future research, including optimizing the time resolution and pooling method for paralinguistic embeddings, using the latest large-scale LLMs, and adding more tasks under serialized multitasking, such as turn taking prediction and emotion recognition.